If y varies inversely as the cube of x, write the particular equation, given that the graph contains (4,7). Find the x if y = 10
Answer by stanbon(75887)

If y varies inversely as the cube of x
---
y = k/x^3
---
write the particular equation, given that the graph contains (4,7).
---
Solve for "k" given that y = 7 when x = 4
7 = k/4^3
k = 7*64 = 448
---
Equation for this Problem:
y = 448/x^3
----
Find the x if y = 10
10 = 448/x^3
x^3 = 448/10
x = 3.5516
Cheers,
Stan H.